# fifth generation 6G sixth generation

### v2

@@
Empirical Validation of a Class of
Ray-Based Fading Models
@@ Abstract
As new wireless standards are developed, the use of higher operation frequencies comes in hand with new use cases and propagation effects that differ from the well-established state of the art. Numerous stochastic fading models have recently emerged under the umbrella of generalized fading conditions, to provide a fine-grain characterization of propagation channels in the mmWave and sub-THz bands. For the first time in literature, this work carries out an experimental validation of a class of such ray-based models, in a wide range of propagation conditions (anechoic, reverberation and indoor) at mmWave bands. We show that the independent fluctuating two-ray (IFTR) model has good capabilities to recreate rather dissimilar environments with high accuracy. We also put forth that the key limitations of the IFTR model arise in the presence of reduced diffuse propagation, and also due to a limited phase variability for the dominant specular components.
@@
channel characterization, stochastic fading channels, IFTR, GTR-V, anechoic chamber, reverberation chamber, indoor environments
00footnotetext: This work has been supported by grant TED2021-129938B-I00 funded by MCIN/AEI/10.13039/501100011033 and by the European Union NextGenerationEU/PRTR. It has also been supported by grants PID2020-112545RB-C54, PDC2022-133900-I00, PDC2023-145862-I00 and TED2021- 131699B-I00, funded by MCIN/AEI/10.13039/501100011033 and by the European Union NextGenerationEU/PRTR; in part by Junta de Andalucía through grant EMERGIA20-00297; and in part by the predoctoral grant FPU22/03392. (Corresponding author: Juan E. Galeote-Cazorla).00footnotetext: Juan E. Galeote-Cazorla, F. Javier Lopez-Martinez and Juan F. Valenzuela-Valdés are with the Dept. Signal Theory, Networking and Communications, Research Centre for Information and Communication Technologies (CITIC-UGR), University of Granada, 18071, Granada, Spain. F. Javier Lopez-Martinez is also with the Communications and Signal Processing Lab, Telecommunication Research Institute (TELMA), Universidad de Málaga, Málaga, 29010, Spain (e-mails: juane@ugr.es; fjlm@ugr.es; juanvalenzuela@ugr.es)00footnotetext: Alejandro Ramírez-Arroyo is with the Department of Electronic Systems, Aalborg University (AAU), 9220 Aalborg, Denmark (e-mail:araar@es.aau.dk).
@@ I Introduction
The advent and widespread deployment of the fifth generation (5G) of wireless technology has been key to improve the achievable data rates well-beyond tenths of Gbps, and also to reduce latency in one order of magnitude with respect to previous standards – now approaching to 1 ms [1]. One of the key enablers for such achievements is the use of higher frequency bands in the millimeter wave (mmWave) range. This trend is likely to continue as 5G evolves, and sixth generation (6G) is often envisioned to provide coverage in the sub-THz range for some use cases [2].
@@
To circumvent the intricacy of these approaches, analytical stochastic channel models are of widespread use due to their comparatively reduced complexity, much fewer parameters, and mathematical tractability [7]. Rayleigh and Rician models are immensely popular in the wireless community, as the de facto standards to model small-scale fading in non line-of-sight (nLoS) and line-of-sight (LoS) cases, respectively. As new use cases promote the ubiquity of wireless devices, many new environments in sensor networks and industrial settings require the development of more advanced models [^1] While any distribution borrowed from statistics may be used to approximate the behavior of wireless channels, only a certain class of distributions that comply with electromagnetic propagation laws will capture the true nature of communications channels., often referred to as generalized. For instance, the two-wave with diffuse power (TWDP) model [8] proposed by Durgin, Rappaport and de Wolf describes a multipath environment with two dominant constant-amplitude specular rays, plus an aggregate diffuse component. With only one parameter addition compared to the Rician case, the TWDP model has an increased capability to recreate not-so-common propagation conditions; these include bimodality and hyper-Rayleigh behavior [9]. Subsequently, different generalizations and alternatives have been proposed in the literature: for instance, the fluctuating two-ray (FTR) model [10] allows that the two rays of the TWDP model fluctuate jointly or independently [11]. This philosophy is inherited from Abdi’s Rician-shadowed model [12], which also generalizes the Rician one. The $\kappa$-$\mu$ and the $\eta$-$\mu$ models proposed by Yacoub add additional flexibility to the Rician and Rayleigh models through the notion of multipath wave clusters. These were also later generalized by Paris’ $\kappa$-$\mu$ shadowed fading model [13].
@@
In this paper, we aim to empirically validate a class of ray-based stochastic fading models in a wide variety of scenarios. Specifically, we used our mixed anechoic/reverberation chamber facilities, together with indoor acquisition, to find a concordance between measurements and different propagation models in the frequency band from 24.25 GHz to 27.5 GHz (i.e., 3GPP n258 band), which belongs to the mmWave general classification [5]. For this purpose, we consider the independent fluctuating two-ray (IFTR) fading model recently proposed in [11], showing a remarkable goodness of fit in most scenarios. Interestingly, we identify that the widespread assumption of uniform phases for the two dominant specular components [8] does not hold in some scenarios with reduced multipath. We show that in those scenarios where this behavior is identified, a more general modeling for the phases [21] improves the fitting performance.
The remainder of the article is organized as follows. Section II summarizes the IFTR propagation model with the relevant physical and mathematical basis. Section III explains how the measurements were acquired and how they are used for recreating new scenarios. Section IV shows and discuss the obtained results from fitting the scenarios with the propagation models. Finally, Section V summarizes the key conclusions and the future lines derived from this work.
@@ II Physical Channel Model
The main advantage of these ray-based statistical models is their physical interpretability. Models such as the mixture of distributions (e.g. Gaussian or Gamma mixture models [25], [26]) can fit experimental data with arbitrary precision. However, they lack any physical interpretation as the number of involved distributions increases. In contrast, models such as the IFTR only need a reduced set of parameters with a clear physical meaning as previously discussed. They allow us to characterize channels in terms of their physical behavior.
III Measurements Analysis
The studied measurement dataset consists in the channel frequency response obtained for three different scenarios, namely: (semi)-anechoic and (semi)-reverberation chambers, and indoor. A more extensive description of the channel environments can be consulted in [27]. They were taken at the Smart and Wireless Applications and Technologies (SWAT) research group facilities, located at the University of Granada (Spain). Acquisition was performed with a vectorial network analyzer (VNA Rohde & Schwarz ZVA67) as the channel sounder, which is capable of measuring the S-parameters up to 67 GHz. The analyzed frequency band sweeps from 24.25 GHz to 27.5 GHz with a 5 MHz step, that is 651 points. It corresponds with the 3GPP n258 band, included in the mmWave classification. Standardized gain horns were used as transmitter (Tx) and receiver (Rx) antennas, specifically the Flann K-band antenna model #21240-20 fed with a WR-34 waveguide-to-coaxial transition. Finally, the transmission power was set to 10 dBm.
As previously mentioned, two of the scenarios were measured in a half anechoic-half reverberation chamber whose dimensions are 5$\times$3.5$\times$3.5 meters (see Fig. 1). The semi-anechoic side is completely covered with absorbents to suppress any possible reflection in order to receive the ray from the LoS path exclusively. In contrast, the semi-reverberation side is composed entirely of metallic walls favoring the appearance of multiple reflections. The emerged behavior corresponds with a multipath channel. In contrast, the indoor scenario is slightly different. It corresponds with the experimental laboratory of the research group, where the chamber is located. In this case, the scenario represents by itself a real environment for communications, whereas the measurements in chambers can be used for the emulation of new channels.
@@ III Measurements Analysis
In the semi-reverberation chamber, the setup changes to a nLoS situation where the Tx antenna is pointing to the reverberation side of the chamber (see Fig. 1). Now, the distance between the antennas is 600 cm, measured as the sum of the distances of each antenna to the metallic wall parallel to the XZ plane. The Tx and Rx antenna configurations are analogous to those described for the anechoic chamber, obtaining additional 1089 channel measurements. For this case, the azimuth angle $\phi$ = 0º represents the situation where the Tx antenna points to the metallic wall parallel to the XZ plane. In this scenario, multipath channels are generated where up to seven reflections are identified in our experiments depending on the Tx-Rx configuration. Since it is a semi-reverberation chamber, the measurements involve several individual rays instead of a continuous reception as one would expect in a full reverberation chamber. This is a consequence of the absorption at the anechoic side. Each ray has a particular power that can be significantly higher than the noise floor or comparable to it, classifying them as specular or diffuse components respectively.
Figure 1: Panoramic picture of the half anechoic-half reverberation chamber. The receiver antenna can move over the hole XZ plane. However, the transmitter only modifies the $\phi$ and $\beta$ angles, named azimuth and roll respectively, for changing orientation and polarization.
@@
Finally, in the indoor setup the Tx antenna is located inside the chamber pointing outside to the Rx antenna in LoS, which is positioned inside the laboratory room (an illustrative figure can be found in [27]). The Rx describes the same virtual URA as for the chamber scenarios. However, the Tx changes the azimuth angle $\phi$, now sweeping -30º, -15º, and 0º for exclusively pointing outside the chamber in a LoS situation. The swept $\beta$ angles have no changes: $\pm$ 30º and 0º. Thus, the total number of measured channels for the indoor scenario is also 1089. As a complex environment, several specular and diffuse components are received with the LoS ray.
@@
A TOSM (Through - Open - Short - Match) calibration process has been performed, setting the calibration planes at the end of the coaxial lines that connect the VNA with the horn antennas. Thus, all the acquired measurements include the antenna gain patterns as components of the propagation channel. Then, we can admit that the channel response $H(f)$ is equivalent to the measured transmission parameter $S_{21}(f)$. On the other hand, the received signal $Y(f)$ from a linear and time-invariant system can be computed as $Y(f)$ = $H(f)\>X(f)$, where $X(f)$ is the transmitted signal. If we assume that $X(f)$ = 1 V, we can admit that all the VNA measurements are expressed in terms of volts. Additionally, each of the 651 frequency points can be considered as an individual realization of the channels because of the static conditions of the environments during the measurement process.
$\hat{H}(f)=H_{1}(f)+H_{2}(f).$ (8)
In the anechoic case a low diffuse power and the presence of only one relevant ray are expected. When two measurements are combined by applying (8), the resulting channel may reproduce scenarios where the diffuse power is reduced and only one or two rays are received in a LoS situation, emulating the two-ray scenario conditions intrinsic to the IFTR model. On the other hand, the reverberation case has a strong diffuse power with the presence of multiple rays. The combination of channels may resemble an industrial environment where nLoS situations are common and reflections over metallic surfaces are the main propagation mechanisms. Additionally, linear combinations may also be performed for the indoor measurements. In this case, the advantage relies on the increase of channel diversity since it is equivalent to introducing new reflections or additional diffuse power sources. The fitting of the IFTR model with these measurements allow us to validate it in realistic scenarios.
@@
The number of possible pairs of linear combinations becomes unmanageable since it is in the order of hundreds of thousands [^3] Two animations have been included as supplementary material for the reader showing numerous configurations examples and their amplitude PDF for the three studied scenarios.. This fact implies that only a reduced sample of all the possible combinations shall be selected [29], searching for a compromise between the number of configurations and the total execution time of the IFTR model validation process. Specifically, 142 different configurations have been randomly chosen (they will be the same ones for the three scenarios). Among them, 131 correspond to cases where the combined pairs share a common Tx orientation, and 11 to pairs sharing a common Rx array position. To avoid possible biases, each sub-set has been previously analyzed. In the first case, it has been checked that the distribution of distances between the positions of the selected pairs is approximately equal to the distribution of all the possible pairs in the Rx array. On the other hand, combinations of different Tx orientations lead to three situations: same polarization but different pointing, same pointing but different polarization, and different pointing and polarization. It has been checked that the random sub-set includes combinations of the three types.
For a fair comparison, the 142 selected configurations are the same for anechoic, reverberation, and indoor scenarios. For the last one, as previously mentioned, the azimuth angle swept -30º, -15º, and 0º instead of -30º, 0º, and 30º. This implies that configurations of anechoic and reverberation scenarios cannot be replicated exactly for the indoor one. However, by tagging the angles as -1, 0, and 1 respectively in both cases, combinations can be computed while keeping the same tag for the three scenarios.
@@
Figure 3: PDFs associated with some examples of merged channels for the three scenarios. Each color represents one individual channel configuration. A great diversity of behaviors is observed (e.g. Rician- and Rayleigh-like, multimodality…).
@@
As a result of performing the linear combinations, a great diversity of channels emerges. Fig. 3 shows some amplitude PDF examples for the three scenarios. A multitude of behaviors are observed: Rician- and Rayleigh-like, multimodality… These latter are especially common in the anechoic configurations [^4] While the merged channels in anechoic case should correspond to a two-ray case by construction, in situations with very reduced multipath the CLT does not apply; hence, additional rays (e.g., those arising from reflections in the posts in Fig. 1) need to be accounted for individually. In this circumstance, the effect of these rays is translated into some additional modality [30].. Bimodality is a particular behavior where two local maxima appear. Ideally, they result from the presence of two rays whose power is substantially higher than the diffuse one or the noise floor [8]. In the case of our experimental measurements, the merged channels are computed to guarantee the presence of (at least) two specular components arriving the receiver.

### v3

@@
Empirical Validation of a Class
of Ray-Based Fading Models
@@ Abstract
As new wireless standards are developed, the use of higher operation frequencies comes in hand with new use cases and propagation effects that differ from the well-established state of the art. Numerous stochastic fading models have recently emerged under the umbrella of generalized fading conditions to provide a fine-grain characterization of propagation channels in the mmWave and sub-THz bands. For the first time in literature, this work carries out an experimental validation of a class of such ray-based models in a wide range of propagation conditions (anechoic, reverberation and indoor scenarios) at mmWave bands. These models allow to characterize the communication channel with a reduced number of physically interpretable parameters. In specific, we show that the independent fluctuating two-ray (IFTR) model has good capabilities to recreate rather dissimilar environments with high accuracy and only four parameters. We also put forth that the key limitations of the IFTR model arise in the presence of reduced diffuse propagation, and also due to a limited phase variability for the dominant specular components.
@@
channel characterization, stochastic fading, IFTR, GTR-V, propagation, anechoic chamber, reverberation chamber, indoor environments
00footnotetext: This work has been supported by grant TED2021-129938B-I00 funded by MCIN/AEI/10.13039/501100011033 and by the European Union NextGenerationEU/PRTR. It has also been supported by grants PID2020-112545RB-C54, PDC2022-133900-I00, PDC2023-145862-I00 and TED2021-131699B-I00, funded by MCIN/AEI/10.13039/501100011033 and by the European Union NextGenerationEU/PRTR; in part by Junta de Andalucía through grant EMERGIA20-00297; and in part by the predoctoral grant FPU22/03392 (Corresponding author: Juan E. Galeote-Cazorla).00footnotetext: Juan E. Galeote-Cazorla, F. Javier Lopez-Martinez and Juan F. Valenzuela-Valdés are with the Department of Signal Theory, Telematics and Communications, Research Centre for Information and Communication Technologies (CITIC-UGR), University of Granada, 18071, Granada, Spain. F. Javier Lopez-Martinez is also with the Communications and Signal Processing Lab, Telecommunication Research Institute (TELMA), Universidad de Málaga, Málaga, 29010, Spain (e-mails: juane@ugr.es; fjlm@ugr.es; juanvalenzuela@ugr.es)00footnotetext: Alejandro Ramírez-Arroyo is with the Department of Electronic Systems, Aalborg University (AAU), 9220 Aalborg, Denmark (e-mail:araar@es.aau.dk).00footnotetext: Digital Object Identifier 10.1109/XXX.2024.XXXXXXX
@@ I Introduction
The advent and widespread deployment of the fifth generation (5G) of wireless technology has been key to improve the achievable data rates well-beyond tenths of Gbps, and also to reduce latency in one order of magnitude with respect to previous standards — now approaching to 1 ms [1]. One of the key enablers for such achievements is the use of higher frequency bands in the millimeter waves (mmWave) range. This trend is likely to continue as 5G evolves, and sixth generation (6G) is often envisioned to provide coverage in the sub-THz range for some use cases [2].
@@
To circumvent the intricacy of these approaches, analytical stochastic channel models are of widespread use due to their comparatively reduced complexity, much fewer parameters, and mathematical tractability [7]. Rayleigh and Rician models are immensely popular in the wireless community, as the de facto standards to model small-scale fading in non line-of-sight (nLoS) and line-of-sight (LoS) cases, respectively. As new use cases promote the ubiquity of wireless devices, many new environments in sensor networks and industrial settings require the development of more advanced models [^1] While any distribution borrowed from statistics may be used to approximate the behavior of wireless channels, only a certain class of distributions that comply with electromagnetic propagation laws will capture the true nature of communications channels., often referred to as generalized. For instance, the two-wave with diffuse power (TWDP) model [8] proposed by Durgin, Rappaport and de Wolf describes a multipath environment with two dominant constant-amplitude specular rays, plus an aggregate diffuse component. With only one parameter addition compared to the Rician case, the TWDP model has an increased capability to recreate not-so-common propagation conditions; these include bimodality and hyper-Rayleigh behavior [9]. Subsequently, different generalizations and alternatives have been proposed in the literature: for instance, the fluctuating two-ray (FTR) model [10] allows that the two rays of the TWDP model fluctuate jointly or independently [11]. This philosophy is inherited from Abdi’s Rician-shadowed model [12], which also generalizes the Rician one. The $\kappa$-$\mu$ and the $\eta$-$\mu$ models proposed by Yacoub add additional flexibility to the Rician and Rayleigh models through the notion of multipath wave clusters. These were also later generalized by Paris’ $\kappa$-$\mu$ shadowed fading model [13]. The goal of ray-based models is to improve the fitting performance in real environments avoiding the aforementioned use of a large number of parameters with unclear physical meaning.
@@
In this paper, we aim to empirically validate [^2] In this context, this terminology is used in a broader sense than in experimental physics literature, where the classical approach for experimental validation is hypothesis testing. However, the state of the art in channel modeling often discards it in favor of a model-based one [20]. That is, empirical validation refers to the process to properly fit and describe measurements with theoretical models previously proposed in the literature. This is the approach we consider here for empirical validation, and the one followed in other works as [21]. a class of ray-based stochastic fading models in a wide variety of scenarios. Specifically, we used our mixed anechoic/reverberation chamber facilities, together with indoor acquisition, to find a concordance between measurements and different propagation models in the frequency band from 24.25 GHz to 27.5 GHz (i.e., 3GPP n258 band), which belongs to the mmWave general classification [22]. For this purpose, we consider the independent fluctuating two-ray (IFTR) fading model recently proposed in [11], which seemed more adequate than the original FTR model [10] or further generalizations of it for this specific context. The IFTR model offers a great flexibility to model the wide spectrum of experimental channels through only four physically interpretable parameters. Finally, this model showed a remarkable goodness of fit (GoF) in most scenarios. Nevertheless, we also identified certain scenarios which were not properly fitted with the IFTR. Interestingly, we identify that the widespread assumption of uniform phases for the two dominant specular components [8] does not hold in some scenarios with reduced multipath. We show that in those scenarios where this behavior is identified, a more general modeling for the phases [23] improves the fitting performance. In brief, we successfully validate these ray-based models in a wide spectrum of environments and rather dissimilar conditions with a detailed analysis of their physically meaningful parameters.
The remainder of the article is organized as follows. Section II summarizes the independent fluctuating two-ray (IFTR) propagation model with the relevant physical and mathematical basis. Section III explains how the measurements were acquired and how they are used for recreating new scenarios. Section IV shows and discuss the obtained results from fitting the scenarios with the propagation models. Finally, Section V summarizes the key conclusions and the future lines derived from this work.
@@ II Physical Channel Model
The main advantage of these ray-based statistical models is their physical interpretability and the reduced number of parameters. Alternatives such as the mixture of distributions (e.g. Gaussian or Gamma mixture models [27][28]) can fit experimental data with arbitrary precision. However, they lack any physical interpretation as the number of involved distributions increases. In contrast, models such as the IFTR only need a reduced set of parameters with a clear physical meaning as previously discussed. They allow us to characterize channels in terms of their physical behavior.
III Measurements Description
The studied measurement dataset consists in the channel frequency response obtained for three different scenarios, namely: semi-anechoic and semi-reverberation chambers, and indoor. A more extensive description of the channel environments can be consulted in [29]. They were acquired at the Smart and Wireless Applications and Technologies (SWAT) research group facilities, located at the University of Granada (Spain). Acquisition was performed with a vector network analyzer (VNA Rohde & Schwarz ZVA67) as the channel sounder, which is capable of measuring the S-parameters up to 67 GHz. The analyzed frequency band sweeps from 24.25 GHz to 27.5 GHz with a 5 MHz step, which leads to 651 measured frequency points. It corresponds with the 3GPP n258 band, included in the mmWave classification. Standardized gain horns were used as transmitter (Tx) and receiver (Rx) antennas, specifically the Flann K-band antenna model #21240-20 fed with a WR-34 waveguide-to-coaxial transition. Finally, the transmission power was set to 10 dBm.
As previously mentioned, two of the scenarios were measured in a semi-anechoic semi-reverberation chamber whose dimensions are 5$\times$3.5$\times$3.5 meters [see Fig. 1(a)]. The semi-anechoic side is completely covered with absorbents to suppress any possible reflection in order to receive the ray from the LoS path exclusively. In contrast, the semi-reverberation side is composed entirely of metallic walls favoring the appearance of multiple reflections. The emerged behavior corresponds with a multipath channel. In contrast, the indoor scenario is slightly different. It corresponds with the experimental laboratory of the research group, where the chamber is located. In this case, the scenario represents by itself a real environment for communications, whereas the measurements in chambers can be used for the emulation of new channels.
(a) Semi-anechoic and semi-reverberation chamber.
(b) Indoor laboratory.
Figure 1: Panoramic picture of (a) the chamber environments, and (b) the indoor scenario. The receiver antenna can move over the entire XZ plane. However, the transmitter only modifies the $\phi$ and $\beta$ angles, named azimuth and roll respectively, for changing orientation and polarization.
@@ III Measurements Description
In the semi-reverberation chamber, the setup changes to a nLoS situation where the Tx antenna is pointing to the reverberation side of the chamber [see Fig. 1(a)]. Now, the distance between the antennas is 600 cm, measured as the sum of the distances of each antenna to the metallic wall parallel to the XZ plane. The Tx and Rx antenna configurations are analogous to those described for the anechoic chamber, obtaining additional 1089 channel measurements. For this case, the azimuth angle $\phi$ = 0º represents the situation where the Tx antenna points to the metallic wall parallel to the XZ plane. In this scenario, multipath channels are generated where up to seven reflections are identified in our experiments depending on the Tx-Rx configuration. Since it is a semi-reverberation chamber, the measurements involve several individual rays instead of a continuous reception as one would expect in a full reverberation chamber. This is a consequence of the absorption at the anechoic side. Each ray has a particular power that can be significantly higher than the noise floor or comparable to it, classifying them as specular or diffuse components respectively.
@@
Finally, in the indoor setup the Tx antenna is located inside the chamber pointing outside to the Rx antenna in LoS, which is positioned inside the laboratory room [see Fig. 1(b)]. It is furnished with several desks, chairs and shelves. The walls are made of concrete and have metal-edge windows. About the setup configuration, the Rx describes the same 11 $\times$ 11 virtual URA as for the chamber scenarios with a separation between elements of 4 cm. However, the Tx changes the azimuth angle $\phi$, now sweeping --30º, --15º, and 0º for exclusively pointing outside the chamber in a LoS situation across the chamber right side window and the door. The swept $\beta$ angles remain the same, that is, $\pm$30º and 0º. Thus, the total number of measured channels for the indoor scenario is also 1089. As a complex environment, several specular and diffuse components are received along with the LoS ray.
@@
A TOSM (Through – Open – Short – Match) calibration process has been performed, setting the calibration planes at the end of the coaxial lines that connect the VNA with the horn antennas. Thus, all the acquired measurements include the antenna gain patterns as components of the propagation channel. Then, we can admit that the channel response $H(f)$ is equivalent to the measured transmission parameter $S_{21}(f)$. On the other hand, the received signal $Y(f)$ from a linear and time-invariant system can be computed as $Y(f)$ = $H(f)\>X(f)$, where $X(f)$ is the transmitted signal. If we assume that $X(f)$ = 1 V, we can admit that all the VNA measurements are expressed in terms of volts. Additionally, each of the 651 frequency points can be considered as an individual realization of the channels because of the static conditions of the environments during the measurement process.
@@
$\hat{H}(f)=H_{1}(f)+H_{2}(f).$ (8)
In the anechoic case a low diffuse power and the presence of only one relevant ray are expected. When two measurements are combined by applying (8), the resulting channel may reproduce scenarios where the diffuse power is reduced and only one or two rays are received in a LoS situation, emulating the two-ray scenario conditions intrinsic to the IFTR model. This is closely related to the appearance of bimodality behaviors in the channel PDFs, an aspect that cannot be described with the classical Rician and Rayleigh models. On the other hand, the reverberation case has a strong diffuse power with the presence of multiple rays. The combination of channels may resemble an industrial environment where nLoS situations are common and reflections over metallic surfaces are the main propagation mechanisms. Additionally, linear combinations may also be performed for the indoor measurements. In this case, the advantage relies on the increase of channel diversity since it is equivalent to introducing new reflections or additional diffuse power sources. The fitting of the IFTR model with these measurements allow us to validate it in realistic scenarios.
@@
The number of possible pairs of linear combinations becomes unmanageable since it is in the order of hundreds of thousands [^4] Two animations have been included as supplementary material for the reader showing numerous configurations examples and their amplitude PDF for the three studied scenarios. They are available in Appendix B.. This fact implies that only a reduced sample of all the possible combinations shall be selected [31], searching for a compromise between the number of configurations and the total execution time of the IFTR model validation process. Specifically, 142 different configurations have been randomly chosen (they will be the same ones for the three scenarios). Among them, 131 correspond to cases where the combined pairs share a common Tx orientation, and 11 to pairs sharing a common Rx array position. To avoid possible biases, each sub-set has been previously analyzed. In the first case, it has been checked that the distribution of distances between the positions of the selected pairs is approximately equal to the distribution of all the possible pairs in the Rx array. On the other hand, combinations of different Tx orientations lead to three situations: same polarization but different pointing, same pointing but different polarization, and different pointing and polarization. It has been checked that the random sub-set includes combinations of the three types. For a fair comparison, the 142 selected configurations are the same for anechoic, reverberation, and indoor scenarios. In this last case, as previously mentioned, the azimuth angle swept --30º, --15º, and 0º instead of --30º, 0º, and 30º. This implies that configurations of anechoic and reverberation scenarios cannot be replicated exactly for the indoor one. However, by tagging the angles as --1, 0, and 1 respectively in both cases, combinations can be computed while keeping the same tag for the three scenarios.
@@
Figure 3: Amplitude PDFs associated with some examples of merged channels for the three scenarios. Each color represents one individual channel configuration. A great diversity of behaviors is observed (e.g. Rician-like and Rayleigh-like, multimodality…).
Figure 4: Cumulative distributions of the obtained IFTR parameters for the three analysed environments.
@@
As a result of performing the linear combinations, a great diversity of channels emerges. Fig. 3 shows some examples of amplitude PDFs for the three scenarios. A multitude of behaviors such as Rician-like, Rayleigh-like or multimodality are observed. These latter are especially common in the anechoic configurations [^5] While the merged channels in anechoic case should correspond to a two-ray case by construction, in situations with very reduced multipath the CLT does not apply; hence, additional rays [e.g., those arising from reflections in the posts, see Fig. 1(a)] need to be accounted for individually. In this circumstance, the effect of these rays is translated into some additional modality [32]., but also monomodal ones. Bimodality is a particular behavior where two local maxima appear. Ideally, they result from the presence of two rays whose power is substantially higher than the diffuse one or the noise floor [8]. In the case of our experimental measurements, the merged channels are computed to guarantee the presence of (at least) two specular components arriving at the receiver.
As mentioned above, a wide spectrum of communication environments are obtained. From a more general point of view, these merged channels might recreate scenarios such as offices, meeting rooms or hallways. The presence of specular and diffuse components would depend on site-specific characteristics (e.g. building materials or obstacles) and the operation frequency, finally resulting in monomodal or multimodal propagation channels. For example, if the LoS component arrives to the receiver together with a strong reflection from a metallic surface, a bimodal channel may naturally emerge.
Frequency is also key to the expected environments behavior. Sub-6 GHz bands propitiate the appearance of a great number of diffuse components due to the basic propagation mechanisms: reflection, diffraction and scattering. This leads naturally to the classical Rician/Rayleigh scenarios depending on the presence/absence of the LoS component [33]. In contrast, for higher frequencies in the sub-THz and THz bands the propagation mechanisms might be different. Diffuse power decreases because of the higher losses and the detriment of some mechanisms such as reflection or diffraction. Then, several individual contributions may stand out from the rest (not only the LoS) naturally emerging multimodal and more complex behaviors [34]. Our particular case, the mmWaves band, is located in between the sub-6 GHz and the THz bands. Depending on the specific characteristic of the propagation scenarios, channels might be closer to the former or the latter case. Therefore, the great diversity of the merged channels exemplified with Fig. 3 becomes an excellent opportunity to validate the ray-based models at the n258 mmWaves band.